Two sides of an equilateral triangle have lengths 3x + 1 and x + 11.
Which of 21 - x or 4x + 5 could be the length of the third side?
a. 21 - x only
b. neither 21 - x nor 4x + 5
c. 4x + 5 only
d. both 21 - x and 4x + 5

Answer by ewatrrr(24785)

 
Hi,
equilateral triangle
3x + 1 = x + 11
x = 5 OR sides are 16
21-5 = 16
4*5 + 5 ≠ 16
a. 21 - x only